how do you solve with slope intercept form and standard form 5y=4x-2 (5, -1)
Answer by joLeeBee(6)

5y = 4x-2
y= 4/5x - 2/5
the intercept form is y = mx + b where m is the slope
in this case your slope is the coefficient of x
so, slope = 4/5